Give a couple of numeral examples to show how coverage ratios help to identify the risk...

  1. Give a couple of numeral examples to show how coverage ratios help to identify the risk or coverage ability on risk. Hint: Times interest earned, EBITDA coverage ratio, cash flow operations to total debt, and free operating cash flow to total debt each have their perspective on the ability of debt paying ability.
  2. Why are liquidity and solvency ratios important to credit analysis? Use the numerical examples to show their utility.
  3. Why is the credit rating necessary for credit analysis? Use one rating system to show the usefulness.
Answer #1

The assignment has more than 1 question and the student has not posted the accurate question requirement, hence, question 2 is solved in totality below:

Why are liquidity and solvency ratios important to credit analysis? Use the numerical examples to show their utility.

ANSWER:

Liquidity ratios asses a company ability to convert its assets to cash and pay off its obligations without any significant difficulty. Liquidity ratios are particularly useful for suppliers, employees, banks, etc. Important liquidity ratios are:

1. Current Ratio

2. Quick Ratio

3. Cash Ratio

Example:

GIVEN:

Current Liabilities 65,000
Current Assets 85,000
Stock 20,000
Advance Tax 5,000
Prepaid Expense 10,000

Quick Ratio =  Quick Assets/Current Liabilities or Quick Liabilities

Quick Assets = All Current Assets – Stock – Prepaid Expenses = 85000 – (20000+5000+10000) = 50,000

Quick Liabilities = All Current Liabilities – Bank Overdraft – Cash Credit = 65,000

Quick Ratio =  50000/65000 = 0.77:1

UTILITY:

This means that for every dollar of Company's current liabilities, the firm has $0.77 of very liquid assets to cover those immediate obligations.

Solvency ratios assess the company's ability to pay off its long-term obligations such as bank loans, bonds payable, etc. Information about solvency is critical for banks, employees, owners, bond holders, institutional investors, government, etc. Key solvency ratios are:

1. Debt ratio

2.Debt-Equity ratio

3. Debt to Capital Ratio

4. Fixed Charge Coverage ratio

5. Times Interest Earned Ratio

6. Equity Multiplier

Example:

GIVEN: NPAT = 1,25,000

Tax Rate = 35%

Debentures are 6,00,000 at 10%

Net Profit before tax = (97500 × 100) ÷ 65

Net Profit Before tax = 1,50,000

Debentures Interest = 6,00,000 × 10% = 60,000

Interest Coverage Ratio = Net Profit before Interest and Tax Interest on Long-Term Debt =  150000/60000

Interest Coverage Ratio = 2.5:1

UTILITY OF RATIO:

We can know, in the current earnings before interest and tax, the firm can cover the interest cost for 2.5 times.